William Michajluk Pauline Partycky m Nicholas Michajluk Mary Michajluk Anna Michajluk Stephen Michajluk 1894-1977 Donna Yurechuk 1905-1991 Dmitro Michajluk Lillian Michajluk m 1927 Ernest George Lake 1928-  Mary Pauline Michajluk 1931-  William Stephen Michajluk 1936-  Ann Elizabeth Dagg m 1952  m 1964 Randall George Lake 1954-  Robert Murray Lake 1957-  Ronald John Willis 1960-  Debra Marilyn Lake 1959- m 1985 Shannon Marie Willis 1988-  Alanna Michelle Willis 1989-  Shawn Douglas Michajluk 1967-  Glen Robertson Lisa Suzanne Michajluk 1970-  Peter William Michajluk 1972- m 1996

1